Give one example for the complex series which is



1) Conditionally Convergent.



2) Absolutely Convergent.



3) Both.



4) Neither.

1
Expert's answer
2022-02-07T15:52:07-0500

1] For conditional convergence;

"\displaystyle
\sum^\infty_{n=1}\frac{(-1)^n}{n}"


2] For absolute convergence;

"\displaystyle
\sum^\infty_{n=1}\frac{i^n}{n^2}"


3] Both absolutely and conditionally convergent; Does not exist since from definition, a complex series that converges but is not absolutely convergent is said to be conditionally convergent. Thus they are exact opposites.


4] For neither;

"\displaystyle
\sum^\infty_{n=0}(1-i)^n"

is a divergent series.